OpenStudy (anonymous):

which expression is the simplified form of sqrt c^15

OpenStudy (anonymous):

\[\sqrt(c^7*c^7*c) = c^7\sqrt(c)\]

OpenStudy (anonymous):

You could also just write:\[\bf c^{\frac{15}{2}}\]Also, I believe there is choices to this question but the asker never provided them. @snewton29 @DoYourHomework